How long will it take $200 to double itself at 9% simple interest
Answer by mananth(16949)

I= p*r*t
p=200
FV = 400
so I = 400-200=200
r=9% = 0.09
t=? years
t= I/p*r
t= 200/ (200*0.09)
=11.11 years